6. (10 points) in the last problem set you graphed the isoquants for using coal (L) or wood (K) to produce heat. One unit of my answers do not feel right to me
Answer #1

Ans.

A) 7.5 C + 5 v = 60

Coal = 60/ 7.5

= 8

Wood = 60/5

= 12

B)

1units of coal produces 5 units of heat

1 unit of wood produces 2 units of heat

Mrts = Pc / Pv

5/2 > 7.5/5

Hence, to maximise the output, he should use more of coal as coal can produce more heat than wood.